# A Comprehensive Characterization of Galaxy-Cool CGM Connections at $z<0.4$ with DESI Year 1 Data

### v1

@@ 3.4 Gas distribution with respect to the size of halos
The results are shown in Figure 10. The Ca II measurements around star-forming and quiescent galaxies are shown in the left and right panels respectively, fitted with
@@
We find that after the normalization, Ca II absorption profiles around star-forming galaxies with different stellar masses become consistent with each other, indicating that the cool gas distributes similarly in the halos, and the behavior of the cool CGM is closely tied to the halo masses of the host galaxies. On the other hand, for quiescent galaxies, there is a different trend, showing that the absorption strengths decrease with stellar masses, especially in the inner regions of the halos $<0.3\,r_{\mathrm{vir}}$. In addition, again, for galaxies with $>10^{10}\,M_{\odot}$, the Ca II absorption around star-forming galaxies is approximately 1.5-2 times stronger than that around quiescent galaxies as observed in the physical space. These results demonstrate that while cool gas traced by Ca II exists in both types of galaxies, they behave differently around different types of galaxies.
@@ 3.5 Ca II mass in the halos
where $f$ is the oscillator strength of Ca II K being 0.648, and $W_{0}^{\mathrm{K}}$ and $\lambda$ are in units of $\rm\AA$. This relation works for unsaturated, optically thin absorption lines and provides a lower limit of column density when the lines are saturated. Our measurements are in the unsaturated region with the strongest $\langle W_{0}^{\mathrm{K}}\rangle$ being $0.2\,\rm\AA$ at $r_{p}>10\,\mathrm{kpc}$ (e.g., Zhu & Ménard, 2013b). We then adopt the best-fit profiles as shown in Figure 10, and integrate the Ca II mass from $0.1r_{\mathrm{vir}}$ to $1r_{\mathrm{vir}}$:
@@
In addition, we estimate the total Ca mass in the ISM of star-forming galaxies by combining the stellar mass-metallicity relation and atomic gas sequence from Scholte et al. (2024) with the solar Ca-to-O abundance ratio of 0.004 (Lodders, 2019). We also apply a dust depletion correction to the metallicity using the relation from Peimbert & Peimbert (2010). The estimated total Ca mass in the ISM of star-forming galaxies is shown by the grey line. For the CGM, we also account for dust depletion in our Ca II measurements of the halos by applying a depletion factor. Using Cloudy simulations (Chatzikos et al., 2023), we determine a depletion factor of $\sim 89$ by comparing the intrinsic Ca II-to-Zn II ratio in the CGM with the observed ratio from Lan & Fukugita (2017). A detailed description of the simulations is provided in Appendix A. After the correction, the amount of the Ca II mass in the halos is comparable to the amount of Ca mass in the ISM. Moreover, the relationship between the Ca II mass and stellar mass of star-forming galaxies follows the relationship of the ISM metal mass and stellar mass of star-forming galaxies. This result demonstrates that the metal enrichment of the ISM and CGM is closely linked together, reflecting the cumulative effect of feedback mechanisms.
Figure 11: Ca II mass in CGM. The solid data points are estimated Ca II mass around star-forming galaxies (blue) and quiescent galaxies (red) and the open data points are values corrected for dust depletion. The grey line is the Ca mass in ISM inferred from the mass-metallicity relation and atomic mass sequence.
@@ 4.1 Comparisons with previous results
By combining DESI spectra of background quasars, we have measured the Ca II absorption strengths around galaxies as a function of stellar mass, galaxy types and redshift, and revealed correlations between Ca II absorption strengths and galaxy properties. In the following, we compare our results with similar measurements from previous studies and discuss the implications.
@@
Cool CGM traced by Ca II at $z<0.2$: Our observed trends are consistent with the trends reported in the previous Ca II results at $z\sim 0.1$ using the SDSS dataset by Zhu & Ménard (2013b). For example, Zhu & Ménard (2013b) separated galaxies into two mass bins and found that galaxies with $10^{10.5}$ and $10^{9.8}\,M_{\odot}$ have best-fit $A_{100}$ parameter values of $15\,\rm m\AA$ and $7\,\rm m\AA$ at 100 kpc. This yields $\beta\sim\rm 0.47$, which is consistent with our best-fit $\beta$ parameter value. We note that while in their analysis, only a mass cut is applied, we expect that the trend in their measurements is mostly driven by signals from star-forming population which has a larger pair number. In addition, they reported that Ca II around star-forming galaxies is approximately two times stronger than Ca II around quiescent galaxies, a trend that we have found in DESI data. We note that limited by the SDSS dataset, Zhu & Ménard (2013b) explored either the dependences between Ca II absorption and mass or SFR, while the DESI dataset enables us to perform the exploration as a function of mass and SFR independently across a wider parameter space than previously explored. We also note that the $\alpha$ parameter of our best-fit profiles is $\sim-2$ for both types of galaxies while Zhu & Ménard (2013b) obtained $\sim-1.4$. The difference is mostly due to the fact that we exclude the $<10\ \,\rm kpc$ measurements during the fitting, while Zhu & Ménard (2013b) included the measurements in their fitting. In Appendix B, we directly compare DESI measurements with SDSS measurements from Zhu & Ménard (2013b), and show that the measurements from the two datasets are consistent.
@@
The only inconsistency between our results and the results in Zhu & Ménard (2013b) is the azimuthal angle dependence. Given that the trend observed in Zhu & Ménard (2013b) is driven by a single measurement in the innermost bin, it is possible that such a trend is affected by small statistics limited by the SDSS data. The larger DESI sample is expected to be less sensitive to the effect. Our measurements are also consistent with the Ca II measurements within 10 kpc of galaxies obtained by Rubin et al. (2022) [^5] We multiply the Ca II K measurements in Rubin et al. (2022) by by a factor of 1.69, which is the weighted mean of the ratio of H+K to K lines from the innermost bins of our measurements. as shown in Figure 3, as well as the Ca II measurements in our Milky Way ISM with typical $W_{0}^{\rm H+K}\sim 0.8-1\rm\,\AA$ from Murga et al. (2015) [^6] Given that the Milky Way sightlines intercept approximately half of the ISM, we multiply the Ca II measurements by two..
@@
Cool CGM traced by Mg II at $z\gtrsim 0.4$: Explorations of the galaxy-cool CGM connections have been conducted with Mg II absorption lines as the tracers of the cool CGM (e.g., Bordoloi et al., 2011; Kacprzak et al., 2012; Rubin et al., 2014; Nielsen et al., 2016; Dutta et al., 2020; Huang et al., 2021), and have revealed similar correlations as found in this work:
@@
  • Mass dependence: Studies have shown that the covering fraction and the absorption profiles of the cool CGM traced by Mg II absorbers increase with stellar mass (and halo mass) of systems. For example, using $\sim 100$ galaxy-quasar pairs, Chen et al. (2010) showed that the characteristic radius of gas traced by Mg II scales with galaxy B-band luminosities as $\propto L_{B}^{0.35}$. With $\sim 200$ galaxy-quasar pairs obtained with MUSE observations, Dutta et al. (2020) also concluded that the stellar mass is the primary factor affecting Mg II absorption around galaxies. Similarly, via a statistical cross-correlation analysis, Lan (2020) demonstrated that the covering fraction of Mg II absorbers measured in physical space correlates with the galaxy mass with $\propto M_{*}^{0.5}$ for star-forming galaxies. Finally, based on cool gas measurements around galaxies and quasars living in halos with different masses, Prochaska et al. (2014) also argued that the halo mass is the primary factor affecting the cool gas distribution, a conclusion which is also reached by Churchill et al. (2013a).
    Together with the Ca II measurements, this mass dependence of the cool CGM has been observed across a wide range of redshifts from $z\sim 0$ up to $z\sim 2$, covering at least 10 Gyr cosmic time. This provides hints on the underlying mechanisms which are expected to operate from $z\sim 2$ and persist through cosmic time.
@@
  • Galaxy type dependence: In addition to the mass dependence, the correlation between galaxy types and the cool CGM, especially in the inner regions, has been reported. Several studies show that the average Mg II absorption strength in the inner CGM around star-forming galaxies is stronger than that around quiescent galaxies (e.g., Bordoloi et al., 2011; Lan et al., 2014; Lan, 2020; Rubin et al., 2018; Anand et al., 2021; Huang et al., 2021). A difference of a factor of $\sim 4$ is observed within 50 kpc, and the value is larger with smaller impact parameters. We note that such a trend is found in Ca II measurements as shown in Section 3. However, the differences in Ca II absorption between star-forming and quiescent galaxies are at most a factor of 2 at $r_{p}\sim 20$ kpc, which is smaller than the results from MgII absorption at higher redshifts. This might suggest that the dependence of the cool CGM on galaxy types becomes weaker at lower redshifts, possibly associated with the declines of the SFR of galaxies and the contribution of galactic outflows. We discuss this implication in the following subsection.
@@ 4.2 Possible redshift evolution of the azimuthal angle dependence
Blue-shifted absorption lines, the observed signature of galactic outflows, have been detected ubiquitously in the spectra of star-forming galaxies, via the so-called down-the-barrel observations (e.g., Weiner et al., 2009; Rubin et al., 2014; Zhu et al., 2015). With the typical observed velocity ($\sim 300$ km/s; e.g., Rubin et al. 2014), gas associated with outflows is expected to travel into the halos and possibly reaches to 50 kpc within a few hundred Myr (e.g., Lan & Mo, 2019). The azimuthal angle dependence of gas distribution has been considered as one of the possible imprints of galactic outflows in the CGM. Such a distribution has been observed in the cool gas traced by Mg II absorption lines. For instance, Bordoloi et al. (2011) measured the Mg II absorption along the minor and major axes of star-forming galaxies at $z\sim 0.7$ with HST ACS images and found that within 50 kpc, the Mg II absorption is stronger along the minor axis than the major axis. The rest equivalent width of Mg II is $\sim 3$ times higher along the minor axis than the major axis at $\sim 25$ kpc. A similar trend and the ratio between Mg II rest equivalent width along the minor and major axes are also found in Lan & Mo (2018) at $z\sim 1$. With individual detections, the bimodal azimuthal dependence of Mg II absorption is reported at $z\sim 0.2$ (e.g., Bouché et al., 2012; Kacprzak et al., 2012). However, the ratio of the Mg II rest equivalent width along the minor to major axes is not as large as the value observed at $z>0.7$ (e.g., Martin et al., 2019; Huang et al., 2021). The azimuthal dependence of Ca II measurements reported in Section 3.3 is more consistent with the $z\sim 0.2$ Mg II results.
@@
Here we propose that the different degree of the azimuthal angle dependence is possibly due to the redshift evolution of galactic outflows driven by star-formation activities. As discussed above, the azimuthal angle dependence is observed mostly at $z>0.7$ (Bordoloi et al., 2011; Lan & Mo, 2018), while the low-redshift CGM has weak or no detectable azimuthal angle dependence as shown in this work and previous studies (e.g., Huang et al., 2021; Borthakur et al., 2015; Pointon et al., 2019). This can be due to the fact that for star-forming galaxies with similar mass, the average SFR at $z\sim 0.8$ is $\sim 3$ times higher than that at $z\sim 0.2$ (e.g., Moustakas et al., 2013). Therefore, three times more gas mass associated with outflows is expected to contribute to the CGM preferentially allocated along the minor axis. Moreover, the velocities of galactic outflows evolve with redshifts. With a fixed SFR, the maximum outflow velocity at $z\sim 1$ is approximately 1.5-2 times faster than the outflow velocity at $z\sim 0$ (e.g., Sugahara et al., 2019). Taking these two factors together, at higher redshifts, galactic outflows are expected to distribute more gas mass into the CGM at larger distances. In other words, the absence of azimuthal angle dependence observed in Ca II at low redshifts indicates that gas associated with galactic outflows does not contribute significantly in the inner CGM. This lack of galactic outflow contribution can also explain why Ca II absorption in the inner CGM around star-forming and quiescent galaxies at $z\sim 0.2$ does not have such a large difference observed at $z>0.7$ with Mg II absorption.
@@
This suppression of galactic outflows towards low redshifts is aligned with the results from the FIRE (Feedback In Realistic Environments) simulations (Hopkins et al., 2014). Muratov et al. (2015) and Anglés-Alcázar et al. (2017) found that high-redshift galaxies drive strong galactic outflows produced by bursty star formation, while at low redshifts, massive galaxies shift into steady star-formation that does not drive strong outflows. Stern et al. (2021) proposed that this transition of the star-formation mode is linked to the virialization of the inner CGM, which occurs at lower redshifts, and after the virialization of the outer CGM. Before the virialization of the inner CGM, the inner CGM is dominated by clumpy cool gas, making galactic outflows relatively easy to travel through. After the virialization of the inner CGM, the hot gas around galaxies confines the galactic outflows. This transition of the inner CGM virialization for $L^{*}$ galaxies occurs at $z\sim 0.8-1$ in the FIRE simulations which coincides with the epoch separating the behaviors of cool gas around star-forming galaxies. While further testing and validation of this outside-in CGM virialization scenario requires direct comparisons between the predicted and observed properties of cool gas tracers, such as Mg II and Ca II, around galaxies across time and mass, the current observed properties of the cool CGM are consistent with the results of the FIRE simulations qualitatively.
@@ 4.3 Impact of AGNs on the properties of the CGM
Our results show that there is no significant difference in the Ca II absorption between AGNs and star-forming galaxies. This indicates that current AGN activities do not have strong impact on the properties of the cool CGM. This result is consistent with previous measurements of gas around galaxies with AGN activities. For example, Berg et al. (2018) selected 19 AGNs via BPT diagnosis and measured the CGM properties with COS UV spectroscopy, finding that the absorption line properties of AGNs are consistent with those of the control sample within 160 kpc. Similarly, Kacprzak et al. (2015) estimated the covering fraction of Mg II absorption around 14 AGNs, and obtained a value slightly lower than that of field galaxies, though within the uncertainties. For luminous AGNs, i.e. quasars, the cool gas distribution has also been measured via quasar-quasar pairs. Previous results show that the properties of cool gas around quasars are similar to those around galaxies with similar stellar mass, suggesting that the halo mass is the primary factor driving the properties of the cool gas (e.g., Farina et al., 2014; Prochaska et al., 2014). On the other hand, Johnson et al. (2015) reported a correlation between the covering fraction of Mg II absorption and quasar luminosity, and argued that quasar feedback can be one of the mechanisms driving the correlation. While validating such a physical connection requires further investigations, it suggests that energy from powerful / luminous quasars may be needed to sufficiently leave imprints on the properties of the cool CGM. Finally, considering radio-mode feedback, Chang et al. (2024) measured the properties of the cool gas traced by Mg II around radio galaxies and the control sample, and found no detectable difference between the properties of the gas around the two populations, illustrating that recent radio AGN activities do not impact the properties of the cool gas strongly. These measurements again support the picture that the properties of galaxies, especially the stellar mass / halo mass and SFR, are the dominant factors correlating with the properties of the cool gas.

### v2

@@ 3.4 Gas distribution with respect to the size of halos
The results are shown in Figure 10. The Ca II measurements around star-forming and quiescent galaxies are shown in the left and right panels, respectively, fitted with
@@
We find that after the normalization, Ca II absorption profiles around star-forming galaxies with different stellar masses become consistent with each other, indicating that the cool gas distributes similarly in the halos, and the behavior of the cool CGM is closely tied to the halo masses of the host galaxies. On the other hand, for quiescent galaxies, there is a different trend, showing that the absorption strengths decrease with stellar masses, especially in the inner regions of the halos $0.3\,r_{\mathrm{vir}}$. In addition, again, for galaxies with $10^{10}\,M_{\odot}$, the Ca II absorption around star-forming galaxies is approximately 1.5–2 times stronger than that around quiescent galaxies as observed in the physical space. These results demonstrate that while cool gas traced by Ca II exists in both types of galaxies, they behave differently around different types of galaxies.
@@ 3.5 Ca II mass in the halos
[…]
@@
In addition, we estimate the total Ca mass in the ISM of star-forming galaxies by combining the stellar mass–metallicity relation and atomic gas sequence from Scholte et al. (2024) with the solar Ca-to-O abundance ratio of 0.004 (Lodders, 2019). We also apply a dust depletion correction to the metallicity using the relation from Peimbert & Peimbert (2010). The estimated total Ca mass in the ISM of star-forming galaxies is shown by the gray line. For the CGM, we also account for dust depletion in our Ca II measurements of the halos by applying a depletion factor. Using Cloudy simulations (Chatzikos et al., 2023), we obtain an estimated depletion factor of $\sim 74$ by comparing the intrinsic Ca II-to-Zn II ratio with the observed ratio in the gas traced by Mg II from Lan & Fukugita (2017) assuming that Ca II and Mg II mostly originate from the same gas in the CGM. This assumption is motivated by the fact that the combination of the Mg II covering fraction (Lan et al., 2014) at $z\sim 0.4$ and the typical Ca II absorption strength in Mg II absorbers (Lan & Fukugita, 2017) yields similar Ca II absorption profiles around galaxies as observed in this work. A detailed description of the simulations is provided in Appendix A. After the correction, the amount of the Ca II mass in the halos is comparable to the amount of Ca mass in the ISM. Moreover, the relationship between the Ca II mass and stellar mass of star-forming galaxies follows the relationship of the ISM metal mass and stellar mass of star-forming galaxies. This result demonstrates that the metal enrichment of the ISM and CGM is closely linked together, reflecting the cumulative effect of feedback mechanisms.
Figure 11: Ca II mass in CGM. The solid data points are estimated Ca II mass around star-forming galaxies (blue) and quiescent galaxies (red) and the open data points are values corrected for dust depletion. The gray line is the Ca mass in ISM inferred from the mass–metallicity relation and atomic mass sequence.
@@ 4.1 Comparisons with previous results
By combining DESI spectra of background quasars, we have measured the Ca II absorption strengths around galaxies as a function of stellar mass, galaxy types, and redshift, and revealed correlations between Ca II absorption strengths and galaxy properties. In the following, we compare our results with similar measurements from previous studies and discuss the implications.
@@
Cool CGM traced by Ca II at $z<0.2$: Our observed trends are consistent with the trends reported in the previous Ca II results at $z\sim 0.1$ using the SDSS data set by Zhu & Ménard (2013b). For example, Zhu & Ménard (2013b) separated galaxies into two mass bins and found that galaxies with $10^{10.5}$ and $10^{9.8}\,M_{\odot}$ have best-fit $A_{100}$ parameter values of $15$ and $7\,\rm m\AA$ at 100 kpc. This yields $\beta\sim\rm 0.47$, which is consistent with our best-fit $\beta$ parameter value. We note that while in their analysis, only a mass cut is applied, we expect that the trend in their measurements is mostly driven by signals from a star-forming population that has a larger pair number. In addition, they reported that Ca II around star-forming galaxies is approximately 2 times stronger than Ca II around quiescent galaxies, a trend that we have found in DESI data. We note that limited by the SDSS data set, Zhu & Ménard (2013b) explored either the dependences between Ca II absorption and mass or SFR, while the DESI data set enables us to perform the exploration as a function of mass and SFR independently across a wider parameter space than previously explored. We also note that the $\alpha$ parameter of our best-fit profiles is $\sim-2$ for both types of galaxies, while Zhu & Ménard (2013b) obtained $\sim-1.4$. The difference is mostly due to the fact that we exclude the $10\ \,\rm kpc$ measurements during the fitting, while Zhu & Ménard (2013b) included the measurements in their fitting. In Appendix B, we directly compare DESI measurements with SDSS measurements from Zhu & Ménard (2013b) and show that the measurements from the two data sets are consistent.
@@
The only inconsistency between our results and the results in Zhu & Ménard (2013b) is the azimuthal angle dependence. Given that the trend observed in Zhu & Ménard (2013b) is driven by a single measurement in the innermost bin, it is possible that such a trend is affected by small statistics limited by the SDSS data. The larger DESI sample is expected to be less sensitive to the effect. Our measurements are also consistent with the Ca II measurements within 10 kpc of galaxies obtained by Rubin et al. (2022) [^5] We multiply the Ca II K measurements in Rubin et al. (2022) by a factor of 1.69, which is the weighted mean of the ratio of H+K to K lines from the innermost bins of our measurements. as shown in Figure 3, as well as the Ca II measurements in our Milky Way ISM with typical $W_{0}^{\rm H+K}\sim 0.8-1\rm\,\AA$ from Murga et al. (2015) [^6] Given that the Milky Way sightlines intercept approximately half of the ISM, we multiply the Ca II measurements by two..
@@
Cool CGM traced by Mg II at $z\gtrsim 0.4$: Explorations of the galaxy–cool CGM connections have been conducted with Mg II absorption lines as the tracers of the cool CGM (e.g., Bordoloi et al., 2011; Kacprzak et al., 2012; Rubin et al., 2014; Nielsen et al., 2016; Dutta et al., 2020; Huang et al., 2021), and have revealed similar correlations as found in this work:
@@
  • Mass dependence: Studies have shown that the covering fraction and the absorption profiles of the cool CGM traced by Mg II absorbers increase with stellar mass (and halo mass) of systems. For example, using $\sim 100$ galaxy–quasar pairs, Chen et al. (2010) showed that the characteristic radius of gas traced by Mg II scales with galaxy $B$-band luminosities as $\propto L_{B}^{0.35}$. With $\sim 200$ galaxy–quasar pairs obtained with MUSE observations, Dutta et al. (2020) also concluded that the stellar mass is the primary factor affecting Mg II absorption around galaxies. Similarly, via a statistical cross-correlation analysis, Lan (2020) demonstrated that the covering fraction of Mg II absorbers measured in physical space correlates with the galaxy mass with $\propto M_{*}^{0.5}$ for star-forming galaxies. Finally, based on cool gas measurements around galaxies and quasars living in halos with different masses, Prochaska et al. (2014) also argued that the halo mass is the primary factor affecting the cool gas distribution, a conclusion that is also reached by Churchill et al. (2013a). Together with the Ca II measurements, this mass dependence of the cool CGM has been observed across a wide range of redshifts from $z\sim 0$ up to $z\sim 2$, covering at least 10 Gyr cosmic time. This provides hints on the underlying mechanisms that are expected to operate from $z\sim 2$ and persist through cosmic time.
@@
  • Galaxy type dependence: In addition to the mass dependence, the correlation between galaxy types and the cool CGM, especially in the inner regions, has been reported. Several studies show that the average Mg II absorption strength in the inner CGM around star-forming galaxies is stronger than that around quiescent galaxies (e.g., Bordoloi et al., 2011; Lan et al., 2014; Rubin et al., 2018; Lan, 2020; Anand et al., 2021; Huang et al., 2021). A difference of a factor of $\sim 4$ is observed within 50 kpc, and the value is larger with smaller impact parameters. We note that such a trend is found in Ca II measurements as shown in Section 3. However, the differences in Ca II absorption between star-forming and quiescent galaxies are at most a factor of 2 at $r_{p}\sim 20$ kpc, which is smaller than the results from Mg II absorption at higher redshifts. This might suggest that the dependence of the cool CGM on galaxy types becomes weaker at lower redshifts, possibly associated with the declines of the SFR of galaxies and the contribution of galactic outflows. We discuss this implication in the following subsection.
@@ 4.2 Possible redshift evolution of the azimuthal angle dependence
[…]
@@
Despite of the abovementioned debate, the current results suggest that the degrees of azimuthal angle dependence vary with different redshifts, being weak or nondetectable at the low-redshift CGM while observed mostly at $z>0.4$ (Bordoloi et al., 2011; Lan & Mo, 2018). We propose that this trend may result from the redshift evolution of galactic outflows driven by star formation activities. For star-forming galaxies with similar mass, the average SFR at $z\sim 0.8$ is $\sim 3$ times higher than that at $z\sim 0.2$ (e.g., Moustakas et al., 2013). Therefore, three times more gas mass associated with outflows is expected to contribute to the CGM preferentially allocated along the minor axis. Moreover, the velocities of galactic outflows evolve with redshifts. With a fixed SFR, the maximum outflow velocity at $z\sim 1$ is approximately 1.5–2 times faster than the outflow velocity at $z\sim 0$ (e.g., Sugahara et al., 2019). Taking these two factors together, at higher redshifts, galactic outflows are expected to distribute more gas mass into the CGM at larger distances. In other words, the absence of azimuthal angle dependence observed in Ca II at low redshifts might indicate that gas associated with galactic outflows does not contribute significantly in the inner CGM. This lack of galactic outflow contribution can also explain why Ca II absorption in the inner CGM around star-forming and quiescent galaxies at $z\sim 0.2$ does not have such a large difference observed at $z\sim 0.7$ with Mg II absorption.
@@
This suppression of galactic outflows toward low redshifts is aligned with the results from the Feedback In Realistic Environments (FIRE) simulations (Hopkins et al., 2014). Muratov et al. (2015) and Anglés-Alcázar et al. (2017) found that high-redshift galaxies drive strong galactic outflows produced by bursty star formation, while at low redshifts, massive galaxies shift into steady star formation that does not drive strong outflows. Stern et al. (2021) proposed that this transition of the star formation mode is linked to the virialization of the inner CGM, which occurs at lower redshifts, and after the virialization of the outer CGM. Before the virialization of the inner CGM, the inner CGM is dominated by clumpy cool gas, making galactic outflows relatively easy to travel through. After the virialization of the inner CGM, the hot gas around galaxies confines the galactic outflows. This transition of the inner CGM virialization for $L^{*}$ galaxies occurs at $z\sim 0.8-1$ in the FIRE simulations, which coincides with the epoch separating the behaviors of cool gas around star-forming galaxies. While further testing and validation of this outside-in CGM virialization scenario requires direct comparisons between the predicted and observed properties of cool gas tracers, such as Mg II and Ca II, around galaxies across time and mass, the current observed properties of the cool CGM are consistent with the results of the FIRE simulations qualitatively.
@@ 4.3 Impact of AGNs on the properties of the CGM
Our results show that there is no significant difference in the Ca II absorption between AGNs and star-forming galaxies. This indicates that current AGN activities do not have strong impact on the properties of the cool CGM. This result is consistent with previous measurements of gas around galaxies with AGN activities. For example, Berg et al. (2018) selected 19 AGNs via BPT diagnosis and measured the CGM properties with COS UV spectroscopy, finding that the absorption-line properties of AGNs are consistent with those of the control sample within 160 kpc. Similarly, Kacprzak et al. (2015) estimated the covering fraction of Mg II absorption around 14 AGNs, and obtained a value slightly lower than that of field galaxies, though within the uncertainties. For luminous AGNs, i.e., quasars, the cool gas distribution has also been measured via quasar–quasar pairs. Previous results show that the properties of cool gas around quasars are similar to those around galaxies with similar stellar mass, suggesting that the halo mass is the primary factor driving the properties of the cool gas (e.g., Farina et al., 2014; Prochaska et al., 2014). On the other hand, Johnson et al. (2015) reported a correlation between the covering fraction of Mg II absorption and quasar luminosity and argued that quasar feedback can be one of the mechanisms driving the correlation. While validating such a physical connection requires further investigations, it suggests that energy from powerful/luminous quasars may be needed to sufficiently leave imprints on the properties of the cool CGM. Finally, considering radio-mode feedback, Chang et al. (2024) measured the properties of the cool gas traced by Mg II around radio galaxies and the control sample and found no detectable difference between the properties of the gas around the two populations, illustrating that recent radio AGN activities do not impact the properties of the cool gas strongly. These measurements again support the picture that the properties of galaxies, especially the stellar/halo mass and SFR, are the dominant factors correlating with the properties of the cool gas.